A falling object has a velocity of 12.3 m/s when it hits a spongy surface. If the object comes to rest in 2.4 cm, how long does it take for the object to come to rest once it has hit the sponge?
Answer #1

acceleration provided by spobge

a = v^2 / 2d = 12.3^2 / (4.8* 10^-2)

a =

using 1st equation of motion

t = v / a =

=======

Comment before rate in case any doubt, will reply for sure.. goodluck